### v2

# Syndrome-Flow Consistency Model Achieves One-step Denoising Error Correction Codes

Abstract
Error Correction Codes (ECC) are fundamental to reliable digital communication, yet designing neural decoders that are both accurate and computationally efficient remains challenging. Recent denoising diffusion decoders achieve state-of-the-art performance, but their iterative sampling limits practicality in low-latency settings. To bridge this gap, consistency models (CMs) offer a potential path to high-fidelity one-step decoding. However, applying CMs to ECC presents a significant challenge: the discrete nature of error correction means the decoding trajectory is highly non-smooth, making it incompatible with a simple continuous timestep parameterization. To address this, we re-parameterize the reverse Probability Flow Ordinary Differential Equation (PF-ODE) by soft-syndrome condition, providing a smooth trajectory of signal corruption. Building on this, we propose the Error Correction Syndrome-Flow Consistency Model (ECCFM), a model-agnostic framework designed specifically for ECC task, ensuring the model learns a smooth trajectory from any noisy signal directly to the original codeword in a single step. Across multiple benchmarks, ECCFM attains lower bit-error-rate (BER) and frame-error-rate (FER) than transformer-based decoders, while delivering inference speeds 30x to 100x faster than iterative denoising diffusion decoders.
1 Introduction
Error Correction Codes (ECC) play a central role in modern digital communications and have been essential in a wide range of applications, including wireless communication and data storage. The core task of an ECC decoder is to recover a message from a received signal corrupted by noise during transmission. Recently, inspired by the great success of deep neural networks, neural network-based decoders were introduced, which are capable of improving the performance scores of conventional, problem-specific algorithms such as Belief Propagation (BP) [33] and Min-Sum (MS) [11].
Figure 1: Illustration of our proposed Error Correction Syndrome-Flow Consistency Model (ECCFM). ECCFM learns to map received signals from the trajectories to a single, consistent codeword prediction $\mathbf{x}_{0}$, represented by a $\delta$ function.
The existing neural decoders can be categorized into two groups. Early model-based approaches [25, 28, 43] achieved successful results by integrating neural networks into conventional decoding algorithms. However, their reliance on problem-specific structures can limit the applicability as a general-purpose decoder. To address this limitation, model-free decoders enable the extension of a general neural network architecture without prior knowledge of decoding algorithms. While the early proposals applying fully connected neural net architectures [13, 3, 17] could lead to overfitting, the preprocessing techniques that utilize magnitude and syndrome vectors [1] have been effective in reducing the error. These improvements have led to efficient transformer-based decoders [5], which leverage the self-attention mechanism to improve the numerical performance on short block codes. The transformer-based decoders have been improved in several recent works [7, 8]. Notably, the recent work by [31] integrates cross-attention between magnitude and syndrome, resulting in the state-of-the-art performance in the decoding task.
In parallel, the remarkable success of diffusion generative models [15, 40] has inspired a new direction for neural decoder. Diffusion models train a noise estimator to gradually denoise noisy input data and reverse a forward noise manipulation process, generating high-quality codewords by an iterative denoising process. DDECC [4] first extends the denoising diffusion framework to ECC, naturally modeling the AWGN channel as a forward diffusion process. A time-dependent transformer learns to denoise the received signal iteratively, recovering the original codeword. This framework provides complementary gains over transformer-based methods, particularly for long codes and low SNR, due to its iterative refinement denoising process.
However, the iterative denoising process inherently suffers from high computational complexity and latency, creating a significant bottleneck for real-time communication systems. While the iterative refinement is the key to high accuracy, it stands in direct conflict with the requirement for fast decoding in practical deployments. This challenge prompts the following question: Can we design a neural decoder that maintains the performance of diffusion-based decoders but with significantly reduced latency?
To solve this tension between accuracy and latency, the consistency model (CM) framework proposed by [37] for general applications provides a potential path. Designed to directly learn the reverse denoising trajectory into a single, efficient step, CMs could potentially offer a direct path to a high-fidelity, one-step decoder. However, our numerical results indicate that a direct application of existing CMs to ECC decoding would not yield the optimal performance, as detailed in Appendix 7.5.7. This outcome appears to arise from a conceptual incompatibility between the continuous nature of CMs and the discrete nature of the ECC task. As we discuss, the core assumption of a standard CM is the existence of a smooth trajectory along a Probability Flow ODE (PF-ODE). In ECC, the level of corruption of a valid codeword is measured by sum of syndrome errors [6]. The decoding path would not be a smooth flow but a series of jumps across a discrete space of error patterns, leading to suboptimal results.
Our main contribution is to resolve this incompatibility by re-parameterizing the decoding trajectory itself. Following the recent work by [20], we introduce a soft-syndrome as a time condition instead of the standard syndrome, which provides a continuous and differentiable measure of the level of corruption of signals. By using the soft-syndrome condition to parameterize the reverse PF-ODE process, we build a smooth trajectory that is native to the ECC problem, mapping the noisy signal towards the manifold of clean codewords. Building on this, we propose the Error Correction Syndrome-Flow Consistency Model (ECCFM), a model-agnostic consistency framework for training single-step decoders. The ECCFM is trained with a well-designed consistency objective that maps any points on the same soft-syndrome-defined trajectory to the original clean codeword, as illustrated in Figure 1. This forces the model to learn the decoding trajectory by a single consistency function.
We numerically evaluate the effectiveness of ECCFM through experiments on a diverse set of standard codes, including BCH, LDPC, Polar, MacKay, and CCSDS. The results show that ECCFM can improve bit-error-rate (BER) and frame-error-rate (FER) compared to leading transformer-based methods, exhibiting strong gains on longer codes with code lengths above $n=200$. Notably, it achieves an inference speedup of 30x to 100x over the existing iterative denoising diffusion methods in a single step, while maintaining comparable decoding performance. In our numerical evaluation, ECCFM offers a model-free training paradigm for building ECC decoders that achieves near state-of-the-art performance with low latency, and requires only a single-step inference suited for real-world communication systems.
2 Related Works
Neural Network-based ECC Decoders. Neural decoders are broadly categorized into model-based and model-free approaches. Model-based methods unfold and enhance conventional algorithms like Belief Propagation (BP) [33] and Min-Sum (MS) [11] with learnable message-passing components. This paradigm, explored across various architectures [9, 18, 19, 25, 28, 29, 26, 43], consistently surpasses its traditional counterparts in performance [27], but can be constrained by the underlying algorithm’s structure and difficulties in capturing long-range dependencies.
In contrast, model-free decoders treat decoding as an end-to-end learning problem. Early fully-connected architectures [13, 3, 17] suffered from overfitting, a challenge later mitigated by pre-processing techniques like magnitude-syndrome decomposition [1]. The advent of Transformers [42] led to ECCT [5], which framed decoding as an auto-regressive sequence-to-sequence task. Subsequent works built upon this foundation to improve generalization (FECCT [7]), enable joint encoder-decoder training (DC-ECCT [8]), and achieve superior performance among auto-regressive methods with cross-attention (CrossMPT [31]). More recently, DDECC [4] introduced a diffusion-based approach, framing decoding as a denoising process and outperforming auto-regressive decoders.
Diffusion Generative Models. Diffusion generative models [35, 15, 40, 16] have achieved state-of-the-art synthesis quality across diverse domains such as images [10, 34, 32], video [14, 2], text [23, 30], and graphs [41, 22, 21]. Their primary limitation, however, is the significant inference latency caused by the iterative denoising process, which has spurred the development of numerous accelerated sampling methods [36, 24].
The diffusion framework is particularly well-suited for ECC, as the AWGN channel provides a forward process [4]. While diffusion-based decoders have set a new state-of-the-art in performance [4, 31], they inherit the same critical issue of high computational cost. To resolve this accuracy-latency trade-off, we turn to Consistency Models (CMs) [37, 38, 12]. CMs are a recent advance in diffusion model acceleration, designed to learn a direct mapping from any point on a denoising trajectory to its origin. This allows for high-fidelity, single-step generation, offering a potential path toward efficient and powerful ECC decoders.
3 Preliminaries
Error Correction Codes. In the error correction code setting, we consider a linear codebook $\mathcal{C}$, defined by a $k\times n$ generator matrix $G$ and an $(n-k)\times n$ parity-check matrix $H$. Note that these matrices satisfy $GH^{\top}=0$ over the binary field $\mathbb{F}_{2}$. The encoder maps a message $m\in\{0,1\}^{k}$ to an $n$-bit codeword $x\in\mathcal{C}\subset\{0,1\}^{n}$ via the linear transformation $x=mG$. The codeword $x$ is then modulated using Binary Phase-Shift Keying (BPSK), where $0\mapsto+1$ and $1\mapsto-1$, resulting in the signal $x_{s}\in\{-1,+1\}^{n}$. We suppose this signal is transmitted over an Additive White Gaussian Noise (AWGN) channel. The received signal $y$ is given by: $y=x_{s}+z$, where the noise vector $z$ is sampled from an isotropic Gaussian distribution, $z\sim\mathcal{N}(0,\sigma^{2}I_{n})$.
The objective of a decoder is to estimate the original codeword $\hat{x}$ from the noisy signal $y$. An essential tool for error detection is the syndrome, calculated from a hard-demodulated formulation of the received signal, $y_{b}$, performed as $y_{b}=\text{bin}(\text{sign}(y))$. Here $\text{sign}(y)$ is +1+1 for $y\geq 0$ and 1-1 otherwise, and the bin function maps $\{-1,+1\}$ back to $\{1,0\}$. The syndrome is then computed as $s=Hy_{b}^{\top}$. An error is detected if the syndrome is a non-zero vector (i.e., $s(y)\neq 0$). Following the pre-processing technique proposed by [1], the input vector to the neural network is $[|y|,s(y)]$ with length n+(nk)n+(n-k) to avoid overfitting.
Consistency Models. CMs [37] built upon diffusion generative models and were introduced to overcome the inference computational costs by enabling fast, one-step generation. The core principle is that: any two points $(\mathbf{x}_{t},t)$ and $(\mathbf{x}_{r},r)$ on the same PF-ODE trajectory should map to the same origin point $\mathbf{x}_{0}$. CMs learn a function $f_{\theta}(\mathbf{x}_{t},t)$ that directly estimates the trajectory from noisy data to clean data with a single step:
$f_{\theta}(\mathbf{x}_{t},t)=\mathbf{x}_{0},$ (1)
The training objective of CMs is to enforce the self-consistency property across a discrete time steps. The continuous time interval $[0,T]$ is discretized into N1N-1 sub-intervals, defined by timesteps $1=t_{1}<\dots<t_{N}=T$. The model is then trained to minimize the following loss, which enforces that the model’s output remains consistent at adjacent points on the same PF-ODE trajectory:
$\mathcal{L}_{\text{Standard-CM}}(\theta)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[w(t)d\left(f_{\theta}(\mathbf{x}_{t},t),f_{\theta}(\mathbf{x}_{r},r)\right)].$ (2)
Here, $f_{\theta}$ is the consistency network being trained and $w(t)$ denotes the time schedule. We provide a detailed review of diffusion generative models and consistency models in the Appendix, Section 7.1.
4 Methodology
4.1 Syndrome-Flow Consistency Property
((a))
Figure 2: Training Dynamics from iterative denoising to 1-step consistency decoding. DDECC’s iterative diffusion denoising learns a noise predictor, $\epsilon_{\theta}(\cdot,e_{t})$, requiring a multi-step iterative process to reverse the noise and decode the codeword. Our ECCFM, $f_{\theta}(\cdot,e^{\dagger})$, directly learns the mapping from any noisy signal to the original clean codeword. By using the smooth soft-syndrome condition ($e^{\dagger}$), it achieves successful decoding in a single step.
While diffusion-based decoders like DDECC [6] achieve high performance, their iterative nature imposes significant computational overhead at inference. To enable fast, single-step decoding, we move beyond iterative refinement and propose a new paradigm based on the Syndrome-Flow Consistency, which defines the behavior of a one-step decoder. The consistency decoder should learn a function, $f_{\theta}$, that deterministically maps any noisy signals, $\mathbf{x}_{t}$, directly to its clean origin, $\mathbf{x}_{0}$, regardless of the noise level $t$. This concept is inspired by the self-consistency principles of Consistency Models (CMs) [37, 38], but is specifically tailored to the deterministic nature of the error correction task. In decoding, the goal is not to sample from a distribution, but to recover a single, correct codeword defined by a Dirac delta function $\delta(\mathbf{x}-\mathbf{x}_{0})$. Formally, we define the Syndrome-Flow Consistency property as follows:
Syndrome-Flow Consistency Property. Given a trajectory $\{y\mathrel{\mathop{\ordinarycolon}}=\mathbf{x}_{t}\}_{t\in[0,T]}$, we learn the consistency function as $f\mathrel{\mathop{\ordinarycolon}}(\mathbf{x}_{t},t;\theta)\mapsto\mathbf{x}_{0}$, holding the Syndrome-Flow Consistency property: for a given ground-truth codeword $\mathbf{x}_{0}$, all points $(\mathbf{x}_{t},t)$ on any trajectory originating from $\mathbf{x}_{0}$ map directly back to it, i.e., $f_{\theta}(\mathbf{x}_{t},t)=\mathbf{x}_{0},\quad\forall t\in[0,T]$. This implies that for any two noisy signals $\mathbf{x}_{t}$ and $\mathbf{x}_{r}$ derived from the same $\mathbf{x}_{0}$, their consistency function outputs must be identical and correct: $f_{\theta}(\mathbf{x}_{t},t)=f_{\theta}(\mathbf{x}_{r},r)=\mathbf{x}_{0},\forall t,r\in[0,T]$. We learn the consistency model such that the conditional distribution is modeled as $p_{\theta}(\mathbf{x}|\mathbf{x}_{t})=\delta(\mathbf{x}-f_{\theta}(\mathbf{x}_{t},t))$, where $f_{\theta}(\mathbf{x}_{t},t)$ directly estimates $\mathbf{x}_{0}$.
A naive application of the standard consistency training objective (Eq. 2) is insufficient, as it only enforces relative consistency by minimizing $d(f_{\theta}(\mathbf{x}_{t},t),f_{\theta}(\mathbf{x}_{r},r))$. This is an indirect and inefficient objection for ECC task, where the ground truth $\mathbf{x}_{0}$ is known during training. We therefore introduce the Error Correction Consistency Loss (EC-CM), which directly enforces our proposed property by minimizing the distance between each predicted outcome and the true codeword $\mathbf{x}_{0}$:
$\mathcal{L}_{\text{EC-CM}}(\theta)$ $\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[w(t)\left[d\left(f_{\theta}(\mathbf{x}_{t},t),\mathbf{x}_{0}\right)+d\left(f_{\theta}(\mathbf{x}_{r},r),\mathbf{x}_{0}\right)\right]],$ (3)
In the ECC domain, we use Binary Cross Entropy (BCE) for the distance measure $d(\cdot,\cdot)$. Our Proposition 4.1 theoretically demonstrates that our optimization objective, $\mathcal{L}_{\text{EC-CM}}$, also serves as an upper bound on the standard consistency loss measured by Total Variation distance. This confirms that our more direct approach inherently enforces self-consistency.
Proposition 4.1.
Let $\mathcal{L}_{\text{Standard-CM}}$ be defined by the Total Variation distance, $\text{TV}(\cdot,\cdot)$, and $\mathcal{L}_{\text{EC-CM}}$ be defined by the Binary Cross Entropy, $\text{BCE}(\cdot,\cdot)$. For any timesteps $t,r$ and ground truth $\mathbf{x}0$, the following semi–triangle inequality holds:
$\text{TV}^{2}\bigl(f_{\theta}(\mathbf{x}_{t},t),f_{\theta}(\mathbf{x}_{r},r)\bigr)$ $\leq\text{BCE}\bigl(f_{\theta}(\mathbf{x}_{t},t),\mathbf{x}_{0}\bigr)+\text{BCE}\bigl(f_{\theta}(\mathbf{x}_{r},r),\mathbf{x}_{0}\bigr).$ (4)
This change in the training objective compels all decoding trajectories to map directly to their origin This fundamental change requires all decoding trajectories mapping to their origin codeword, as demonstrated in Figure 1, and the learned solution distribution is expected to center on $\mathbf{x}_{0}$.
4.2 Soft-syndrome Condition for Smooth Trajectory
Figure 3: Decoding trajectories for models conditioned on Hard Syndrome versus Soft Syndrome on a POLAR(64,48) code. The soft-syndrome conditioning results in a smoother path to a valid codeword. Additional Results under low SNR are available in Appendix 7.5.2, Figure 7.
The training objective of a consistency model is rooted in the differential equation $\frac{\mathrm{d}f}{\mathrm{d}t}=0$. Following Eq. 10 and 11 in [12], the consistency function $f_{\theta}$ is parameterized to satisfy these conditions:
$f_{\theta}(\mathbf{x}_{t},t)=\mathbf{x}_{0}\Leftrightarrow\frac{\mathrm{d}f}{\mathrm{d}t}=0,f_{\theta}(\mathbf{x}_{0},0)=\mathbf{x}_{0},$ (5)
In practice, this differential form is discretized for training using a finite-difference approximation, by dividing the time horizon into N1N-1 sub-intervals $1=t_{1}<\cdot\cdot\cdot<t_{N}=T$:
$0=\frac{\mathrm{d}f}{\mathrm{d}t}\approx\frac{f_{\theta}(\mathbf{x}_{t},t)-f_{\theta}(\mathbf{x}_{r},r)}{t-r},$ (6)
where $\mathrm{d}t=\Delta t=t-r$, $t>r\geq 0$.
A critical problem arises when applying the consistency framework to decoding tasks: the time variables $t$ and $r$, which measure noise levels, are not directly observable from the received signals. A natural solution, proposed in DDECC [6], is to use the sum of syndrome error, $e_{t}$, as a measurement of the noise level, since $e_{t}=0$ indicates a valid codeword. The hard syndrome is computed as $s=Hy_{b}^{\top}$, where $H$ is the parity-check matrix. The sum of syndrome error, $e_{t}=\sum_{i}s(\mathbf{x}_{t})_{i}$, is then the sum of binary syndrome bits.
However, replacing the continuous time variables $t$ and $r$ in Eq. 6 with their discrete, integer-valued counterparts $e_{t}$ and $e_{r}$ violates the core assumption of a smooth, differential trajectory in consistency training. As we demonstrate in Figure 3, the trajectory defined by the syndrome error is highly non-smooth and clustered. A small change in the noisy signal can cause an abrupt jump in the syndrome error count, invalidating the finite-difference approximation and leading to instability during training in 3.
Therefore, the discrete and non-smooth nature of the syndrome error makes it an unsuitable conditioning variable for consistency training in ECC, which necessitates a smooth and differential measure of noise level along the denoising trajectory. We then propose replacing this discrete error sum condition with a continuous and differentiable alternative. Inspired by [20], we introduce the soft syndrome as the basis for the consistency noise condition. The soft syndrome, $s^{\dagger}$, is a fully differentiable function that leverages the log-likelihood ratios of the received signal $\mathbf{x}_{t}$: and offers a continuous measure of how close each parity-check equation is to being fulfilled.
Algorithm 1 Syndrome-Flow Consistency Training
1:Model $f_{\bm{\theta}}$, parity-check matrix $\mathbf{H}$, learning rate $\eta$, syndrome weight $\lambda$, denoising steps NN, time scaling factor $\alpha$.
2:for training batch $\mathbf{x}_{0}$ do
3:  $t\sim\mathcal{U}\{1,\dots,N\},\,r=\alpha t$ $\triangleright$ Sample forward diffusion steps
4:  $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ $\triangleright$ Sample Gaussian noise
5:  $\mathbf{x}_{t}\leftarrow\mathbf{x}_{0}+\sqrt{\bar{\beta}_{t}}\cdot\bm{\epsilon},\mathbf{x}_{r}\leftarrow\mathbf{x}_{0}+\sqrt{\bar{\beta}_{r}}\cdot\bm{\epsilon}$ $\triangleright$ Generate noisy signals via forward process
6:  $e^{\dagger}_{t}=\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{t},H),\,\,e^{\dagger}_{r}=\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{r},H)$ $\triangleright$ Calculate soft-syndrome
7:  $\mathcal{L}_{\text{EC-CM}}\leftarrow\text{BCE}(f_{\theta}(\mathbf{x}_{t},e^{\dagger}_{t}),\mathbf{x}_{\text{0}})+\text{BCE}(f_{\theta}(\mathbf{x}_{r},e^{\dagger}_{r}),\mathbf{x}_{\text{0}})$
8:  $\mathcal{L}_{\text{Total}}\leftarrow\mathcal{L}_{\text{Consistency}}+\lambda\cdot\left(\mathcal{L}_{\text{Soft-syn}}(f_{\theta}(\mathbf{x}_{t},e^{\dagger}_{t}),H)+\mathcal{L}_{\text{Soft-syn}}(f_{\theta}(\mathbf{x}_{r},e^{\dagger}_{r}),H)\right)$
9:  $\bm{\theta}\leftarrow\bm{\theta}-\eta\cdot\nabla_{\bm{\theta}}\mathcal{L}_{\text{Total}}$
10:end for
Differential Soft-syndrome Error Time Condition. Similar to the conventional hard sum of syndrome error, we compute the soft-syndrome error condition $e^{\dagger}_{t}$ of the parity-check matrix $H$ as:
$e^{\dagger}_{t}$ $\mathrel{\mathop{\ordinarycolon}}=-\frac{1}{n-k}\sum_{j}\log\Pr(S_{j}^{\dagger}=0)=-\frac{1}{n-k}\sum_{j}\log(1-s_{j}^{\dagger}),$ (7)
This soft-syndrome error condition is computed as the binary cross-entropy between the estimated syndrome and the all-zero syndrome. We use the mean-field approximation to provide differential conditions to estimate the probability of satisfying zero-syndrome conditions in Eq. 7:
$s_{j}^{\dagger}=\frac{1}{2}-\frac{1}{2}$ $\prod_{\{i\mathrel{\mathop{\ordinarycolon}}H_{j,i}=1\}}\Bigl(2\cdot\mathrm{sigmoid}\bigl(\frac{2\mathbf{x}_{t,i}}{\sigma^{2}}\bigr)-1\Bigr),$ (8)
$\Pr(S_{j}^{\dagger}=0)=1-s_{j}^{\dagger},$
where $x_{t,i}$ is the $i$-th position of the received codeword $x_{t}$, the noise level $\sigma$ is given to the decoder at the receiver’s side. The soft-syndrome is zero if and only if the codeword is valid, yet it varies smoothly and continuously with the received signal $\mathbf{x}_{t}$. By using $e^{\dagger}_{t}$ as the time condition, we provide the consistency model with a smooth trajectory from a noisy signal to a valid codeword, resolving the instability and degeneracy issues of the hard error sum and enabling stable training.
4.3 Syndrome-Flow Consistency Training Dynamics
[…]
$\mathcal{L}_{\text{Total}}(\theta)=\mathbb{E}\Bigl[\underbrace{w(t)\left(d(f_{\theta}(\mathbf{x}_{t},e^{\dagger}_{t}),\mathbf{x}_{0})+d\left(f_{\theta}(\mathbf{x}_{r},e^{\dagger}_{r}),\mathbf{x}_{0}\right)\right)}_{\text{Consistency Loss}}$ (9)
$+\lambda\cdot\underbrace{\left(\mathcal{L}_{\text{Soft-syn}}(f_{\theta}(\mathbf{x}_{t},e^{\dagger}_{t}),H)+\mathcal{L}_{\text{Soft-syn}}(f_{\theta}(\mathbf{x}_{r},e^{\dagger}_{r}),H)\right)}_{\text{Soft-syndrome Loss}}\Bigr]$
where $\mathcal{L}_{\text{Soft-syn}}(\mathbf{x}_{t},H)=-e^{\dagger}_{t}$, $d(\cdot,\cdot)$ is a distance metric, such as Binary Cross-Entropy (BCE) in this work, and $\lambda$ is a hyperparameter that weights the syndrome regularization term. Once trained according to Algorithm 1, the learned consistency function $f_{\theta}$ can decode the noisy received signal $y$ in a one step: $\hat{\mathbf{x}}_{0}=f_{\theta}(y,e^{\dagger}_{t})$, in Appendix 2, Algorithm 2.
5 Numerical Results
Datasets. We evaluate our proposed ECCFM framework on the following set of standard error correction codes, including BCH, Polar, and LDPC codes (MacKay, CCSDS, and WRAN variants). Our evaluation considers multiple code lengths ($n$), rates ($k/n$), and Signal-to-Noise Ratios (SNRs), specifically $E_{b}/N_{0}$ values from 4 to 6 dB, to ensure a robust assessment of performance.
Evaluation Metrics. We evaluate decoding performance using two standard metrics following established benchmarks [5, 6, 31]: Bit Error Rate (BER) and Frame Error Rate (FER). BER measures the fraction of individual bits that are incorrectly decoded. FER measures the fraction of entire codewords that contain one or more bit errors. Concerning the latency, we evaluate computational efficiency by reporting inference time and throughput (decoded samples per second).
Baselines. We numerically compared the results with multiple baselines for the decoding task, including: 1) Conventional BP-based decoders: BP [1] and ARBP [29]. 2) Transformer-based decoders: ECCT [5], and CrossMPT [31]. 3) Denoising diffusion decoders: DDECC [6].
Experimental Setup. We reproduced the results for all model-free baselines (ECCT, CrossMPT, DDECC) by implementing them with their publication-stated hyperparameters. Our primary ECCFM model utilizes a Cross-attention Transformer following CrossMPT, using $N=6$ layers and a hidden dimension of $d=128$. All the other three baselines (ECCT, CrossMPT, DDECC) apply the same network architecture to ensure fair comparison. ECCFM was trained for 1500 epochs using the Adam optimizer on a single GPU. The learning rate was managed by a cosine decay scheduler, starting at $10^{-4}$ and decreasing to $5\times 10^{-7}$. Detailed training configurations and hyperparameter selections are provided in Appendix 7.4.
Overall Performance. Following established benchmarks [5, 6, 31], we conducted a decoding performance comparison measured in $\ln(\text{BER})$. Our method was evaluated versus two classes of decoders: conventional BP-based algorithms (BP and ARBP) and model-free neural decoders (ECCT, CrossMPT, and DDECC). To ensure a fair comparison, all neural models were implemented with a fixed architecture ($N=6$ layers, $d=128$ hidden dimensions). Furthermore, to ensure statistical significance, each simulation was run until at least 500 error codes were observed, under a maximum of $10^{8}$ test instances. As shown in Table 1, our proposed ECCFM framework consistently achieves the best or second-best BER across the tested code families, including BCH, Polar, LDPC, CCSDS and MacKay with different code rates $(n,k)$, and showing considerable gain over POLAR codes.
Table 1: Performance comparison of various decoders across different codes and Signal-to-Noise Ratios ($E_{b}/N_{0}$). The results are reported in terms of $\ln(\text{BER})$ (the higher, the better). All model-free methods use a fixed model architecture ($N=6$, $d=128$). Best results are shown in bold and the second-best results are shown in underline, respectively.
Architecture BP-based decoders Model-free decoders
Code Type Parameters BP ARBP ECCT CrossMPT DDECC ECCFM(Ours)
4 5 6 4 5 6 4 5 6 4 5 6 4 5 6 4 5 6
BCH (63,36) 4.03 5.42 7.26 4.57 6.39 8.92 4.69 6.48 9.06 4.94 6.74 9.28 5.02 6.82 9.88 5.00 6.89 9.76
(63,45) 4.36 5.55 7.26 4.97 6.90 9.41 5.47 7.56 10.51 5.73 7.98 10.80 5.68 8.08 11.22 5.70 8.03 11.04
POLAR (64,32) 4.26 5.38 6.50 5.57 7.43 9.82 6.87 9.21 12.15 7.42 9.94 13.28 7.04 9.44 12.70 7.55 10.31 13.80
(64,48) 4.74 5.94 7.42 5.41 7.19 9.30 6.21 8.31 10.85 6.36 8.53 11.09 5.93 8.00 10.44 6.56 8.78 11.52
(128,64) 4.10 5.11 6.15 4.84 6.78 9.30 5.79 8.45 11.10 7.45 9.71 14.31 7.71 11.40 13.85 8.01 12.22 16.71
(128,86) 4.49 5.65 6.97 5.39 7.37 10.13 6.29 8.98 12.82 7.43 10.80 15.13 7.61 10.50 13.88 7.78 11.21 16.05
(128,96) 4.61 5.79 7.08 5.27 7.44 10.20 6.30 9.04 12.40 7.06 10.25 13.23 7.14 10.31 13.66 7.21 10.52 14.32
LDPC (121,60) 4.82 7.21 10.87 5.22 8.31 13.07 5.12 8.21 12.80 5.75 9.42 15.21 5.42 9.11 13.82 6.02 9.94 15.55
(121,70) 5.88 8.76 13.04 6.45 10.01 14.77 6.30 10.11 15.50 7.06 11.29 17.10 6.91 11.02 17.15 7.35 12.23 17.60
(121,80) 6.66 9.82 13.98 7.22 11.03 15.90 7.27 11.21 17.02 7.87 12.65 17.72 7.61 11.89 16.18 8.25 13.33 18.69
MacKay (96,48) 6.84 9.40 12.57 7.43 10.65 14.65 7.37 10.55 14.72 7.85 11.72 15.49 8.03 12.44 15.79 7.92 12.25 16.08
CCSDS (128,64) 6.55 9.65 13.78 7.25 10.99 16.36 6.82 10.60 15.87 7.56 11.87 16.80 7.77 12.35 17.22 7.95 12.68 17.01
Performance on Longer Codes. To further evaluate scalability, we conducted a evaluation on longer codes commonly used in practical communication systems: LDPC(n=204,k=102)(n=204,k=102), LDPC(n=529,k=440)(n=529,k=440), WRAN(n=384,k=320)(n=384,k=320), and Polar(n=512,k=384)(n=512,k=384). All methods were implemented with the same model architecture ($N=6,d=128$). As illustrated in Figure 4, ECCFM improves the performance across a range of SNRs (2 dB to 6 dB). Additional results on other high-length codes are presented in Appendix 7.5.8, which also suggest inference speed gain compared to the baselines.
Figure 4: Performance comparison of various decoding baselines on medium-to-long block codes. The plot shows the Bit Error Rate (BER) at different Signal-to-Noise Ratios (SNRs), from 2 dB to 6 dB and divided by 0.5 dB.
Performance on Rayleigh Fading Channel. While our proposed ECCFM training dynamics and the definition of soft-syndrome are intrinsic to the code structure and independent of the underlying channel model, we validate our approach on Rayleigh fading channels to demonstrate robustness. We follow the same experimental settings described in [5, 31]. In contrast to the AWGN channel, the received signal is defined by: $y=hx+z$, where $h$ is an $n$-dimensional i.i.d. Rayleigh distributed vector with a scale parameter $\alpha=1$ and $z\sim N(0,\sigma^{2})$. In Figure 6, our method achieved competitive performance compared to baselines across different SNRs on Polar and LDPC without modifications.
Figure 5: Comparison of Inference Time (top) and Throughput (bottom) across various decoding baselines and code types.
Figure 6: Performance comparison of various decoding baselines on Rayleigh Fading channels. The plot shows the Bit Error Rate (BER) at different Signal-to-Noise Ratios (SNRs).
Inference Time and Throughput Comparison. To quantify the efficiency gain, we measured inference time (total seconds to decode $10^{5}$ samples) and throughput (samples decoded per second). As shown in Figure 6, ECCFM demonstrates a speed advantage over diffusion-based methods such as DDECC, achieving speedups of over 30x for short codes and 100x for medium-to-long codes. This disparity arises because diffusion models require several iterative denoising steps for inference, a computational cost that scales with code complexity as detailed in Appendix 7.5.9. Therefore, ECCFM matches the competitive performance of denoising diffusion decoders while operating at the high throughput of single-step transformer-based decoders.
Ablation Study: Model-Agnostic Property of ECCFM. We discussed that ECCFM is a model-agnostic training framework, i.e., its performance could be preserved over different neural network architectures. We conducted an ablation study where we decoupled our framework from the backbone. Specifically, we used the ECCT baseline architecture and trained it with our proposed ECCFM and compared this model against the original ECCT and DDECC. The results in Appendix 7.5.4, Table 5 show that applying the ECCFM training objective yields improvement in $\ln(\text{BER})$ over the standard ECCT, with comparable performance versus the iterative denoising DDECC method.
Ablation Study: Necessity of Soft-syndrome Time Condition in Consistency Training. We further investigated the necessity of using soft-syndrome as a time condition for building the smooth trajectory. We trained ECCFM with either a hard-syndrome time condition or a soft-syndrome time condition. In Appendix 7.5.6, Table 7, our experiments revealed that using soft-syndrome as time condition is necessary for successful training of ECCFM. The continuous nature of the soft-syndrome provides a smooth trajectory to learn, whereas models conditioned on the hard syndrome failed to converge.
6 Conclusions and Limitations
In this work, we introduced the Error Correction Syndrome-Flow Consistency Flow Model (ECCFM), a novel training framework for obtaining high performance with low latency required for real applications. By reformulating the decoding task as a one-step consistency mapping and introducing a differential soft-syndrome condition, we handle non-smooth trajectories that previously hindered the application of consistency models to ECC. Our experiments demonstrate that ECCFM achieves comparable BER and FER across various standard codes, while offering a consistent inference speedup over denoising diffusion methods. Despite these successful results, our work has limitations to be addressed in future research. The convergence rate and training efficiency of consistency models highly rely on building a smooth trajectory, which varies in different code types. Future work can explore adaptive methods for more general decoding tasks.